Appendix 1: Experiment to find the enthalpy change for the reaction


Zn(s) + Cu2+(aq) → Cu(s) + Zn2+(aq)



2. Pipette 25 cm3 of the copper sulphate solution into a polystyrene cup.

3. Weigh about 6 g of zinc powder in the weighing bottle. (since this is in excess

there is no need to be accurate).

4. Put the thermometer through the hole in the lid,stir and record the

temperature to the nearest 0.1 oC every half minute for 2.5 minutes.

5. At precisely three minutes,add the zinc powder to the cup. 


Calculate ∆H for quantities used Calculate ∆H for 1 mol of each Calculate % error

Comparison w/ accepted value


1
Expert's answer
2021-03-31T07:57:13-0400

25g of Zn was added to 25cm³ of 1M CuSO4 solution in a polystyrene cup over 8 minutes. The temperature of the solution was observed to decrease from 68°C to 25°C.

Moles of Zn= 25/65.4= 0.38mol

Molar mass of CuSO4= 160g/mol

Mole of CuSO4= c x V = 1 x 25/1000= 0.025mol

Mass of CuSO4= mole x Molar mass= 0.025x160= 4g

q= mc∆T= 4 x 4.2 x (25-68)= -722.4J

This is the heat lost by CuSO4.

Since the temperature decreases, heat is negative and process is exothermic

∆H= q/n= -722.4/0.025= -28896J/mol